# Emergent Social Dynamics of LLM Agents in the El Farol Bar Problem

### v2

@@ 3.1 Observation
Finally, in one representative run, we observed that in the later stages of the simulation agents began to communicate using hashtags such as “#collaboration” and “#positivity” (Figure 6). These hashtags spread rapidly within subgroups and sometimes across the entire population, suggesting the spontaneous emergence of collective norms. However, this phenomenon did not occur in every trial, and its statistical robustness will be addressed in the analysis section.
Figure 6: Emergence and spread of hashtags in messages after 900 steps. The background color represents agent groups. In this snapshot, “#collaboration” and “#positivity” first appeared in the orange group and then spread rapidly across other groups, eventually becoming global trends. At this stage, synchronous sharing of hashtags did not appear to correlate with synchronous movement or gathering, but rather with the content of exchanged messages.
@@ 4 Discussion
Nevertheless, the agents did not fully optimize attendance, remaining in an uncomfortable state above the threshold. From the conversations, it is clear that social motivations often outweighed rational optimization, with agents clustering together even before the bar became crowded (Figure 9). In some cases, groups even coordinated voluntary exits after discussions (see Appendix 1). Whether the imperfect solution reflects bounded rationality [21], and whether larger or more precise LLMs would converge to a more optimal equilibrium, remains an open question.

### v3

@@ 3.1 Observation
Notably, in one representative run, we observed that in the later stages of the simulation agents began to communicate using hashtags such as “#collaboration” and “#positivity” (Figure 6). These hashtags spread rapidly within subgroups and sometimes across the entire population, suggesting the spontaneous emergence of collective norms. While a simple visual inspection did not reveal an immediate link to coordinated movement, the functional role of this emergent communication will be quantitatively investigated in the subsequent analysis section.
Figure 6: Emergence and spread of hashtags in messages after 900 steps. The background color represents agent groups. In this snapshot, “#collaboration” and “#positivity” first appeared in the orange group and then spread rapidly across other groups, eventually becoming global trends.
@@ 3.2 Analysis
To understand the factors influencing an agent’s decision to either stay in or leave the crowded bar, we conducted a detailed analysis of their behavior leading up to the overcrowding event. We first identified the time step ($T_{over}$) at which the bar first became crowded (attendance $\geq$ 60%) in each simulation run. Agents inside the bar at $T_{over}$ were then classified into two groups based on their subsequent actions: the “leave group,” who exited the bar within 50 steps, and the “stay group,” who remained. We then calculated the continuous duration each agent had spent inside the bar immediately prior to $T_{over}$ and compared it between the groups. The stay group exhibited a significantly longer prior duration in the bar compared to the leave group (Figure 12). A Welch’s t-test confirmed that this difference is statistically significant ($p<0.01$), indicating that an agent’s decision to remain in a crowded environment is strongly correlated with the amount of time it has already invested in being there.
Figure 12: Comparison of continuous steps spent in the bar prior to overcrowding ($T_{over}$). Agents who remained in the bar after it became crowded (stay group) had spent significantly more time inside beforehand compared to those who left (leave group). The difference is statistically significant (Welch’s t-test, $p<0.01$).
Furthermore, to investigate the dynamic impact of hashtags on group retention, we conducted an event-centered aggregate analysis. For each simulation, the time step of the first hashtag appearance was defined as the temporal reference point (step 0). We then aligned all simulation trials around this reference point and calculated the average exit rate (number of leavers / total agents in the bar) over time (Figure 13). A notable feature of this trend is that the sharp decline in the exit rate appears to begin several steps prior to the hashtag’s appearance at step 0. Comparing the periods before and after the event, the exit rate dropped sharply and subsequently tended to remain at a low level for several steps. This dramatic shift quantitatively demonstrates that the emergence of hashtag communication plays a critical role in temporarily suppressing agent departure from the group.
Figure 13: Average agent exit rate centered on the first appearance of a hashtag (step 0), aggregated across 10 simulation runs. The solid line represents the mean exit rate, and the shaded area indicates the standard deviation. The exit rate from the bar drops dramatically immediately following the appearance of a hashtag and remains at a low level for several steps.
@@ 4 Discussion
Nevertheless, the agents did not fully optimize attendance, remaining in an uncomfortable state above the threshold. Our analysis reveals interconnected mechanisms behind this imperfect optimization. First, agents exhibited a form of behavioral inertia: those who had already invested more time in the bar were significantly less likely to leave once it became crowded (Figure 12). This tendency, where past actions weigh on present choices, may prevent a purely rational exit. Second, the use of hashtags emerged as a key social mechanism reinforcing this inertia. Interestingly, the decline in exit rate slightly precedes the appearance of the hashtag (Figure 13), suggesting that the hashtag itself is not the direct trigger for the behavioral shift, but rather serves to crystallize a collective sentiment or agreement that was already forming. This suggests that shared cultural symbols, like hashtags, function as an emergent norm that strengthens group cohesion and encourages agents to resist physical discomforts like crowding. These findings, combined with the general observation that social motivations often outweighed rational optimization (Figure 9), indicate that agent behavior is driven by an interplay between individual history and collective cultural formation. In some cases, groups even coordinated voluntary exits after discussions (see Appendix 1). Whether this imperfect solution reflects bounded rationality [21], and whether larger or more precise LLMs would converge to a more optimal equilibrium, remains an open question.